A 4.0-g bullet leaves the muzzle of a rifle with a speed of 340 m/s. What force (assumed constant) is exerted on the bullet while it is traveling down the 0.75-m-long barrel of the rifle?


1
Expert's answer
2022-06-14T08:42:11-0400

Given:

"m=0.004\:\rm kg"

"v_0=0,\quad v=340\:\rm m/s"

"d=0.75\:\rm m"


The energy-work theorem says

"\frac{mv^2}{2}-\frac{mv_0^2}{2}=Fd"

Thus, the force exerted onj the bullet

"F=\frac{mv^2}{2d}=\frac{0.004*340^2}{2*0.75}=310\:\rm N"